OpenStudy (anonymous):

-4(x+5)^2 +11 = -245 Solve quadratic function with square roots

Nnesha (nnesha):

do you understand how to solve these type of questions ? thats the same as last one what would be the first step ?

OpenStudy (retireed):

Simplify it into the form ax^2 + bx + c = 0

Nnesha (nnesha):

by mentions this `with square roots` i don't think they wants em to foil (x+5)^2 you can take square root both sides to cancel out the square but in order to do that you need to get rid of everything except (x+5)^2 at left side

OpenStudy (retireed):

okay this is new to me "with square roots" so let's subtract 11 from both sides

OpenStudy (retireed):

-4(x+5)^2 =256 right then take the square root of both sides

OpenStudy (danjs):

recall a square root is the same as raised to 1/2 power \[\sqrt{a}=a^{1/2}\]

OpenStudy (danjs):

first isolate what is being squared by itself, by dividing both sides by (-4)

OpenStudy (danjs):

(x+5)^2 = 64

OpenStudy (retireed):

-2(x+5) =16 x+5 = -8 x = -13 this might be wrong so follow DanJS 's path

OpenStudy (danjs):

then remember for a square root of something squared, results in an absolute value \[\sqrt{(x+5)^2}= \left| x+5 \right|\]

OpenStudy (danjs):

so you end up with \[\left| x+5 \right|=8\]

OpenStudy (danjs):

to solve those , the absolute value term can equal + or - 8 x+5=+8 or x+5=-8 x has 2 values